Reduce the second order linear differential equation d²y/dt²-7dy/dt+10y=0 to linear system of first order differential equation and hence solve the system of ODE's


1
Expert's answer
2021-06-15T18:26:08-0400
"\dfrac{d^2y}{dt^2}-7\dfrac{dy}{dt}+10y=0"




"y''-7y'+10y=0"

Let "u=y, v=y'(t)."

Then "u'=y'=v, v'=y''=-10y+7y'=-10u+7v"



"u'=0u+1v""v'=-10u+7v"


"\begin{bmatrix}
 u' \\
 v'
\end{bmatrix}=
\begin{bmatrix}
 0 & 1 \\
 -10 & 7
\end{bmatrix}
\begin{bmatrix}
 u \\
 v
\end{bmatrix}"

"\vec {x'}=\begin{bmatrix}
 u' \\
 v'
\end{bmatrix}, A=\begin{bmatrix}
 0 & 1 \\
 -10 & 7
\end{bmatrix}, \vec x=\begin{bmatrix}
 u \\
 v
\end{bmatrix}"

The linear system of first order differential equation 

"\vec {x'}= A\vec x"

Find the eigenvalues and eigen vectors of "A=\begin{bmatrix}
 0 & 1 \\
 -10 & 7
\end{bmatrix}"



"\det (A-\lambda I)=A=\begin{vmatrix}
 0-\lambda & 1 \\
 -10 & 7-\lambda
\end{vmatrix}"

"=-\lambda(7-\lambda)+10=\lambda^2-7\lambda+10"

"\det (A-\lambda I)=0=>\lambda^2-7\lambda+10=0"

"\lambda_1=5, \lambda_2=2"

These are the eigenvalues.

"\lambda=5"


"\begin{bmatrix}
 0-\lambda & 1 \\
 -10 & 7-\lambda
\end{bmatrix}=\begin{bmatrix}
 -5 & 1 \\
 -10 & 2
\end{bmatrix}"

"R_2=R_2-2R_1"


"\begin{bmatrix}
 -5 & 1 \\
 0 & 0
\end{bmatrix}"

"R_1=-R_1\/5"


"\begin{bmatrix}
 1 & -1/5 \\
 0 & 0
\end{bmatrix}"

If we take "q_2=1," then "q_1=1\/5."


"\vec q=\begin{bmatrix}
 1/5 \\
 1
\end{bmatrix}"

"\lambda=2"

"\begin{bmatrix}
 0-\lambda & 1 \\
 -10 & 7-\lambda
\end{bmatrix}=\begin{bmatrix}
 -2 & 1 \\
 -10 & 5
\end{bmatrix}"

"R_2=R_2-5R_1"

"\begin{bmatrix}
 -2 & 1 \\
 0 & 0
\end{bmatrix}"

"R_1=-R_1\/2"

"\begin{bmatrix}
 1 & -1/2 \\
 0 & 0
\end{bmatrix}"

If we take "s_2=1," then "s_1=1\/2."

"\vec s=\begin{bmatrix}
 1/2 \\
 1
\end{bmatrix}"


General solution is


"\begin{bmatrix}
 u \\
 v
\end{bmatrix}=C_1e^{5t}\begin{bmatrix}
 1 \\
 5
\end{bmatrix}+C_2e^{2t}\begin{bmatrix}
 1 \\
 2
\end{bmatrix}"


"y=C_1e^{5t}+C_2e^{2t}"